# Politecnico di Torino, DIMEAS, Corso Duca degli Abruzzi, 24 10129 Torino, Italy

### v1

@@ II.2 Qubits and Quantum States
To be self-contained, we provide a brief outline of the key elements of quantum computers. A detailed introduction can be found in Nielsen and Chuang (2011). In quantum computers, the internal state of the machine can be represented by a complex vector in a Hilbert space. The time evolution during computation is described by a matrix multiplication of this vector with a unitary operator.
In most quantum hardware architectures, the elementary building blocks, known as qubits, are microscopic systems that can be described by a complex-valued, normalized, two-dimensional Hilbert vector in which the computational basis states $|0\rangle$ and $|1\rangle$ (kets) can be associated with the Cartesian basis vectors, and the complex-valued amplitudes $\alpha$ and $\beta$ are normalized as $|\alpha|^{2}+|\beta|^{2}=1$
@@
Most quantum states (Eq. (6)) are entangled and therefore cannot be decomposed into such a product. Unitary evolution amounts to a matrix multiplication, i.e. the final quantum state after a computation is
@@ II.3 Hybrid Quantum-Classical Reservoir Computing
Hybrid quantum-classical reservoir computing Pfeffer et al. (2022) is based on principles that are similar to classical reservoir computing introduced in Sec. II.1. In QRC, the exponential size of $n$-qubit Hilbert spaces is utilized to encode and process classical reservoir state vectors, $\boldsymbol{r}(t_{i})$. Thus, the classical reservoir is replaced by a quantum counterpart. More specifically, within a single loop of the training phase, the reservoir state $\boldsymbol{r}(t_{i})$ is mapped onto an $n$-qubit unitary operator $P(\boldsymbol{r}(t_{i}))$, that takes $|0\rangle^{\otimes n}$ to the state $|\psi(t_{i})\rangle=P(\boldsymbol{r}(t_{i}))|0\rangle^{\otimes n}$. The explicit form of $P$ is given by the quantum circuit architecture that characterizes the specific QRC framework. Circuits are generally parametrized by single-qubit rotation angles and a number of entangling CNOT gates (Appendix A).
@@
For a training data set $\{\boldsymbol{\hat{u}}(t_{0}),\boldsymbol{\hat{u}}(t_{1}),\ldots,\boldsymbol{\hat{u}}(t_{N_{tr}})\}$, the input time-series is encoded in the quantum circuit using single-qubit rotation angles. With this encoding, a second unitary operator $\Phi(\boldsymbol{u_{in}}(t_{i}))$ is applied to the quantum state $|\psi(t_{i})\rangle$. Thereafter, a third unitary operator $V(\alpha)$ given by a random parameterized circuit with $n$ parameters $\alpha$ is applied. The combined action for each time step is in Fig. 5
${|\psi(t_{i+1})\rangle}=V(\alpha)\,\Phi(\boldsymbol{u}_{in}(t_{i}))\,P(\boldsymbol{r}(t_{i}))|0\rangle^{\otimes n}.$ (9)
@@
Similarly to classical reservoir computing, we derive a pre-processed reservoir state vector $\hat{\boldsymbol{r}}(t_{i+1})$ from measured probabilities of $|\psi(t_{i+1})\rangle$ in the computational basis. This requires sampling and measuring the state $|\psi(t_{i+1})\rangle$ multiple times for each time stepping.
Following the post-processing of $\hat{\boldsymbol{r}}(t_{i+1})$ to $\boldsymbol{r}(t_{i+1})$ (Eq. (2)), the reservoir state matrix $\boldsymbol{R}$ is formed by concatenating the reservoir state vectors. As in classical reservoir computing, the initial reservoir state is chosen randomly, which in turn, necessitates the washout of a few initial time-stepped reservoir vectors.
In the training phase, QRC is run in open loop, and the output matrix $\boldsymbol{W}_{out}$ is computed by a linear ridge regression (Eq. (3)). When operated in a closed loop, QRC is used to predict the system dynamics (Eq. (4)). We compare the elementary steps of classical and quantum reservoir computing algorithms in Fig. 3. The open-loop and closed-loop configurations of the hybrid quantum-classical architecture are presented in Fig. 4. Each reservoir update $\boldsymbol{r}({t}_{i})$ in both the open-loop and closed-loop requires the input of the previous reservoir state $\boldsymbol{r}({t}_{i-1})$, similarly to the classical reservoir update (dotted feedback loop in Fig. 4). In Sec. III, we propose a quantum-classical architecture, which is independent of this feedback loop, thereby allowing for parallelisation of the training phase of QRC.
@@
Figure 4: Quantum-classical reservoir architecture schematic representation (a) open-loop training to produce $\boldsymbol{W}_{out}$ matrix (b) closed-loop autonomous predictions starting from an arbitrary point in the unseen data set. The dotted line in both figures indicates the recurrence involved in conventional reservoir architectures. In RF-QRC architecture, we remove this recurrent feedback layer of reservoir states.
Figure 5: Schematic representation of gate-based quantum reservoir computer Pfeffer et al. (2022). The first two unitaries are used to encode previous reservoir states $\boldsymbol{r}(t_{i})$ and input time-series $\boldsymbol{u}_{in}(t_{i})$, the third additional unitary provide randomization. Each unitary consists of a specific feature map. The quantum circuit execution starts from the ground state $|\boldsymbol{0}\rangle$ and measures the reservoir state activation for the next time-step $\hat{\boldsymbol{r}}(t_{i+1})$, which is post-processed classical before feeding the next time-step. Each reservoir activation state $\boldsymbol{r}(t_{i+1})$ is saved classically to use for training and predictions.
@@
Although QRC exploits the exponential size of the state space of the quantum register, there is a practical limitation due to the limited connectivity of the employed quantum processors (e.g. neighbour-neighbour interactions only on 2D superconducting qubit processors). The performance of QRC generally depends on the degree of the entanglement present in the circuit Götting et al. (2023). Therefore, we investigate the performance of structurally different architectures.
@@ III Recurrence-Free Quantum Reservoir Computing (RF-QRC):
The gate-based quantum reservoir computing framework of Pfeffer et al. (2022) (Sec. II.3) is in principle equivalent to its classical counterpart. It has been used to study thermal convection flows and the turbulent Rayleigh Bernard flow Pfeffer et al. (2022, 2023). In this section, we propose new ansätze for the circuit architecture in QRC. To do so, a specific quantum feature map must be chosen and we therefore start by briefly reviewing a number of common choices. Following that, we present the Recurrence-Free Quantum Reservoir Computing architecture, which will be employed to study low and high-dimensional chaotic flows in Sec. IV.
We consider four different feature maps for QRC. These are the (a) Linearly Entangling feature map, used in Pfeffer et al. (2022) for QRC applications, (b) Product State feature map, which was proposed in Havlíček et al. (2019), and has been shown to offer a high expressivity Abbas et al. (2021), (c) Fully Entangling feature map, in which CNOT gates entangle pairs of qubits, (d) Symmetric Fully Entangling feature map, which only differs from (c) by an additional data encoding layer following the CNOT gates. These feature maps do not require anywhere-to-anywhere connectivity (which would increase the computational overhead on current superconducting quantum processors). Schematic circuit diagrams of all four feature maps are shown in Appendix A. When the number $\Theta$ of encoded parameters of a string of data (e.g. for $\boldsymbol{r}$, $\Theta=2^{n}$), exceeds the number of qubits, the feature map encoding is applied multiple times. In particular, this leads to exponentially growing circuit depths of the unitary $P$ and motivates the exploration of new efficient QRC architectures, as displayed in Tab. 1.
@@
QRC-C4 - Fully entangled (x2) Fully entangled symmetric
@@
An example comparison of circuit depths corresponding to the five architectures is shown in Fig. 6. Removing the recurrence, i.e. by setting $P$ to the identity, and by applying $\Phi$ twice for stimulating the reservoir dynamics, yield circuits depths that are independent of the reservoir size. The suitability of this choice is justified a posteriori by an improved performance of QRC in a range of prototypical model systems. By removing the recurrence, QRC becomes similar to Quantum Extreme Learning Machines (QELM) Mujal et al. (2021). However, the difference lies in a linear combination of the memory and nonlinearity of previous reservoir activation states as expressed by Eq. (2) (also known as the leaky-integral reservoir computing approach Jaeger et al. (2007)).
@@ IV Numerical Results:
In this section, we analyse the dynamics of a number of chaotic systems, such as Lorenz-63, Lorenz-96, and MFE (Moehlis, Faisst, and Eckhardt) using CRC and QRC. Thereby, we can assess the prediction capabilities of various architectures. More specifically, we evaluate the short-term time-accurate, and long-term statistical predictions with performance measures. The sets of training, validation, and test data are generated by a $4^{th}$-order Runge-Kutta numerical scheme. To emulate the quantum circuits, we have used the Qiskit Qiskit contributors (2023) software package to compute the noise-free evolution of the quantum register.
@@ IV.1 Performance Measures
The dynamics of chaotic systems can be characterized by the leading Lyapunov Exponent $\Lambda_{1}$ Boffetta et al. (2002). This exponent corresponds to the average exponential rate of divergence for initially nearby trajectories. The Lyapunov Exponent also provides a time scale to assess the time-accurate prediction of the chaotic systems. We have re-scaled our time units to the inverse of the Lyapunov Exponent ($\Lambda_{1}$), which is called the Lyapunov Time ($1LT=\Lambda_{1}^{-1}$).
In this work, we choose the Valid Prediction Time (VPT) Vlachas et al. (2020) as a performance measure of short-term time-accurate predictions. For a given threshold value $\varepsilon$, VPT is defined as the time for which the Normalized Root Mean Square Error (NRMSE) between the predicted and true values is less than $\varepsilon$. In this work we take $\varepsilon$ = 0.5, as in Vlachas et al. (2020)
$NRMSE=\sqrt{||{\frac{\boldsymbol{y}_{t}-\boldsymbol{\hat{y}}_{t}}{\sigma^{2}}}||},$ (10)
@@
where $\boldsymbol{y}_{t}$ is the true value and $\boldsymbol{\hat{y}}_{t}$ is the corresponding predicted value, and $\sigma$ is the standard deviation of the time series, $t_{f}$ is the largest value of the time step at which NRMSE is smaller than the threshold $\varepsilon$. High VPT values indicate a high predictability. To assess the long-term statistical prediction capabilities of our reservoir networks, we evolve the trained networks autonomously in a closed-loop configuration from an arbitrary point on the attractor and compare the state variable distributions with the actual distributions. For each configuration, we run simulations with five different sets of optimal hyperparameters, which corresponds to different random seeds.
@@ IV.2 Three-dimensional Lorenz-63 model
As a first test case, we start with the analysis of the three-dimensional Lorenz-63 system Lorenz (1995). The Lorenz-63 system is a reduced-order model of a thermal convection flow, in which the fluid is heated uniformly from below and cooled from the top. This model is defined mathematically by
@@
where [$\sigma$ , $\rho$ , $\beta$] = [10, 28, 8/3] results in a chaotic behavior of the system. For the Lorenz-63 system, $\Lambda=0.9$ and $0.91LT=1/$ 0.9, Lorenz (1995). The time series data set is derived by a Runge-Kutta method for $dt=0.01s$. The training time series comprises data points over a total time of 20 LT for both the classical and quantum reservoir networks. For the training, both networks are executed in an open loop to evolve the reservoir states and calculate the $\boldsymbol{W}_{out}$ matrix. For predictions, both classical and quantum networks evolve dynamically in the closed-loop configuration from ensembles of points in phase space, sampled on the attractor randomly, to quantify the time accuracy. These points correspond to the points in the unseen test data set.
In Tab. 2, we compare the hyperparameters of both quantum and classical reservoirs. Values inside square brackets indicate that the parameters are optimized within this range, multiple values indicate that we repeated the hyperparameter search for each of these values and then selected the best VPT out of them. The random rotation angles ($\alpha$) in the third unitary block can also be treated as a hyperparameter. Alternatively, one could derive them from a uniform random distribution, which we shall do in this paper.
@@ IV.2.1 Time-accurate predictions
Figure 7: Lorenz-63 system, closed-loop time-series predictions. Comparison of true predictions with classical reservoir computing (CRC) and emulated quantum reservoir computing (QRC).
Figure 8: Lorenz-63 system. VPT/circuit depth vs. reservoir sizes for both classical and quantum reservoir configurations.
Figure 9: Lorenz-63 system. VPT vs. reservoir sizes for different quantum reservoir configurations.
In Fig. 8, we present the VPT values derived with classical reservoir computing and compare them with different configurations of quantum reservoir computers from Tab. 1. The error bars at each value indicate the variation arising from different random seeds. We have tested five different seeds for each value. These seeds correspond to randomly generated weight matrices $\boldsymbol{W}$ and $\boldsymbol{W}_{in}$ for the classical reservoir, and variation of $\alpha$ values for quantum reservoir architectures. For each of these choices, after training, we initialize the closed-loop predictions from 20 different initial points for quantifying VPT and its mean value.
@@
The prediction performance of the two Recurrence-Free QRC configurations, QRC-C3 and QRC-C5, do not improve with the reservoir sizes. The QRC-C4 architecture, however, shows comparable performance to the reservoirs with feedback loops (QRC-C1 and QRC-C2). These results highlight the importance of carefully chosen feature maps for QRC.
In Fig. 9, the VPT values normalized by the circuit depth are compared between different QRC architectures. We find that the proposed QRC-C4 architecture outperforms all other architectures by demonstrating the best prediction capabilities whilst requiring a smaller circuit depth.

### v2

@@ II.2 Qubits and quantum states
To be self-contained, we provide a brief outline of the key elements of quantum computers. A detailed introduction can be found in Ref. [44]. In quantum computers, the internal state of the machine can be represented by a complex vector in a Hilbert space. The time evolution during computation is described by a matrix multiplication of this vector with a unitary operator.
In most quantum hardware architectures, the elementary building blocks, known as qubits, are microscopic systems that can be described by a complex-valued, normalized, two-dimensional vector in a Hilbert space, in which the computational basis states $|0\rangle$ and $|1\rangle$ (kets) can be associated with the Cartesian basis vectors, and the complex-valued amplitudes $\alpha$ and $\beta$ are normalized as $|\alpha|^{2}+|\beta|^{2}=1$
@@
Most quantum states [Eq. (6)] are entangled and therefore cannot be decomposed into such a product. Unitary evolution amounts to a matrix multiplication, i.e. the final quantum state after a computation is
@@ II.3 Hybrid quantum-classical reservoir computing
Hybrid quantum-classical reservoir computing [30] is based on principles that are similar to classical reservoir computing introduced in Sec. II.1. In QRC, the exponential size of $n$-qubit Hilbert spaces is utilized to encode and process classical reservoir state vectors, $\bm{r}(t_{i})$. Thus, the classical reservoir is replaced by a quantum counterpart. More specifically, within a single loop of the training phase, the reservoir state $\bm{r}(t_{i})$ is mapped onto an $n$-qubit unitary operator $P(\bm{r}(t_{i}))$, that takes $|0\rangle^{\otimes n}$ to the state $|\psi(t_{i})\rangle=P(\bm{r}(t_{i}))|0\rangle^{\otimes n}$. The explicit form of $P$ is given by the quantum circuit architecture that characterizes the specific QRC framework. Circuits are generally parametrized by single-qubit rotation angles and a number of entangling CNOT gates (Appendix A).
@@
For a training data set $\{\bm{\hat{u}}(t_{0}),\bm{\hat{u}}(t_{1}),\ldots,\bm{\hat{u}}(t_{N_{tr}})\}$, the input time-series is encoded in the quantum circuit using single-qubit rotation angles. With this encoding, a second unitary operator $\Phi(\bm{u_{in}}(t_{i}))$ is applied to the quantum state $|\psi(t_{i})\rangle$. Thereafter, a third unitary operator $V(\alpha)$ given by a random parameterized circuit with $n$ parameters $\alpha$ is applied. The combined action for each time step is in Fig. 6
${|\psi(t_{i+1})\rangle}=V(\alpha)\,\Phi(\pmb{u}_{in}(t_{i}))\,P(\pmb{r}(t_{i}))|0\rangle^{\otimes n}.$ (9)
@@
Similarly to classical reservoir computing, we derive a pre-processed reservoir state vector $\hat{\bm{r}}(t_{i+1})$ from measured probabilities of $|\psi(t_{i+1})\rangle$ in the computational basis. This requires sampling and measuring the state $|\psi(t_{i+1})\rangle$ multiple times for each time stepping.
[…]
@@
Figure 5: Quantum-classical reservoir architecture schematic representation (a) open-loop training to produce $\pmb{W}_{out}$ matrix (b) closed-loop autonomous predictions starting from an arbitrary point in the unseen data set. The dotted line in both figures indicates the recurrence involved in conventional reservoir architectures. In RF-QRC architecture, we remove this recurrent feedback layer of reservoir states.
In the training phase, QRC is run in an open-loop and the output matrix $\bm{W}_{out}$ is computed by a linear ridge regression [Eq. (3)]. When operated in a closed-loop, QRC is used to predict the system dynamics [Eq. (4)]. We compare the elementary steps of classical and quantum reservoir computing algorithms in Fig. 4. The open-loop and closed-loop configurations of the hybrid quantum-classical architecture are presented in Fig. 5. Each reservoir update $\pmb{r}({t}_{i})$ in both the open-loop and closed-loop requires the input of the previous reservoir state $\pmb{r}({t}_{i-1})$, similarly to the classical reservoir update (dotted feedback loop in Fig. 5). In Sec. III, we propose a quantum-classical architecture, which is independent of this feedback loop, thereby allowing for parallelization of the training phase of QRC.
[…]
@@
Although QRC exploits the exponential size of the state space of the quantum register, there is a practical limitation due to the limited connectivity of the employed quantum processors (e.g. neighbour-neighbour interactions only on 2D superconducting qubit processors). The performance of QRC generally depends on the degree of the entanglement present in the circuit [45]. Therefore, we investigate the performance of structurally different architectures.
@@ III Recurrence-Free Quantum Reservoir Computing (RF-QRC):
QRC-C4 (RF-QRC) - Fully entangled (x2) Fully entangled symmetric
@@
The gate-based quantum reservoir computing framework of [30] (Sec. II.3) is in principle equivalent to its classical counterpart. It has been used to study thermal convection flows and the turbulent Rayleigh-Bernard flow [30, 46]. In this section, we propose an ansätze for the circuit architecture in QRC. To do so, a specific quantum feature map must be chosen and we therefore start by briefly reviewing a number of common choices. Following that, we present the RF-QRC architecture, which will be employed to study low and high-dimensional chaotic flows in Sec. IV.
We consider four different feature maps for QRC. These are the (a) Linearly entangling feature map, used in Ref. [30] for QRC applications, (b) Product state feature map, which was proposed in Ref. [47], and has been shown to offer a high expressivity [48], (c) Fully entangling feature map, in which CNOT gates entangle pairs of qubits, (d) Symmetric fully entangling feature map, which only differs from (c) by an additional data encoding layer following the CNOT gates. These feature maps do not require anywhere-to-anywhere connectivity (which would increase the computational overhead on current superconducting quantum processors). Schematic circuit diagrams of all four feature maps are shown in Appendix A. When the number $\Theta$ of encoded parameters of a string of data (e.g. for $\bm{r}$, $\Theta=2^{n}$), exceeds the number of qubits, the feature map encoding is applied multiple times. In particular, this leads to exponentially growing circuit depths of the unitary $P$ and motivates the exploration of new efficient QRC architectures, as displayed in Tab. 1.
@@
An example comparison of circuit depths corresponding to the five architectures is shown in Fig. 7. Removing the recurrence, i.e. by setting $P$ to the identity, and by applying $\Phi$ twice for stimulating the reservoir dynamics, yield circuits depths that are independent of the reservoir size. The suitability of this choice is justified a posteriori by an improved performance of QRC in a range of prototypical model systems. Recurrence-free QRC is different than Quantum Extreme Learning Machines (QELMs) [49], because of the use of a linear combination of the memory and nonlinearity of previous reservoir activation states, as expressed by Eq. (2) (also known as the leaky-integral reservoir computing approach [50]).
@@ IV Numerical Results:
In this section, we analyze the dynamics of a number of chaotic systems, such as Lorenz-63, Lorenz-96, and MFE using CRC and QRC. Thereby, we can assess the prediction capabilities of various architectures. More specifically, we evaluate the short-term time-accurate and long-term statistical predictions with performance measures. The sets of training, validation, and test data are generated by a $4^{th}$-order Runge-Kutta numerical scheme. To emulate the quantum circuits, we have used the Qiskit [51] software package to compute the noise-free evolution of the quantum register.
In a physical implementation, at each time step between $10^{4}$ to $10^{5}$ shots are required to reconstruct the reservoir activation signal. A number of shots between $10^{4}$ to $10^{6}$ is common in variational algorithms and other QML applications [52]. Mitigating the requirement of a large number of shots in quantum computing is an area of active research, which is beyond the scope of this paper. (QRC and RF-QRC could be combined with the concept of weak and projective measurements [52], classical shadows [53], Resolvable Expressive Capacity (REC) [54], or artificial memory restriction [55] to potentially reduce the number of required shots.)
@@ IV.1 Performance measures
The dynamics of chaotic systems can be characterized by the leading Lyapunov exponent $\Lambda_{1}$ [38]. This exponent corresponds to the average exponential rate of divergence for initially nearby trajectories. The Lyapunov exponent also provides a time scale to assess the time-accurate prediction of the chaotic systems. We have re-scaled our time units to the inverse of the Lyapunov exponent ($\Lambda_{1}$), which is called the Lyapunov Time ($1LT=\Lambda_{1}^{-1}$).
In this work, we choose the Valid Prediction Time (VPT) [9] as a performance measure of short-term time-accurate predictions. For a given threshold value $\varepsilon$, VPT is defined as the time for which the Normalized Root Mean Square Error (NRMSE) between the predicted and true values is less than $\varepsilon$. In this work we take $\varepsilon$ = 0.5, as in Ref. [9]
$NRMSE=\sqrt{||{\frac{\pmb{y}_{t}-\pmb{\hat{y}}_{t}}{\sigma^{2}}}||},$ (10)
@@
where $\pmb{y}_{t}$ is the true value and $\pmb{\hat{y}}_{t}$ is the corresponding predicted value, and $\sigma$ is the standard deviation of the time series, $t_{f}$ is the largest value of the time step at which NRMSE is smaller than the threshold $\varepsilon$. High VPT values indicate a high predictability. To assess the long-term statistical prediction capabilities of our reservoir networks, we evolve the trained networks autonomously in a closed-loop configuration from an arbitrary point on the attractor and compare the state variable distributions with the actual distributions. For each configuration, we run simulations with five different sets of optimal hyperparameters, which corresponds to different random seeds.
@@ IV.2 Three-dimensional Lorenz-63 model
As a first test case, we start with the analysis of the three-dimensional Lorenz-63 system [36]. The Lorenz-63 system is a reduced-order model of a thermal convection flow, in which the fluid is heated uniformly from below and cooled from the top. This model is defined mathematically by
@@
where [$\sigma$ , $\rho$ , $\beta$] = [10, 28, 8/3] results in a chaotic behavior of the system. For the Lorenz-63 system, $\Lambda=0.9$ and $0.91LT=1/$ 0.9, [36]. The time series data set is derived by a Runge-Kutta method for $dt=0.01s$. The training time series comprises data points over a total time of 20 LT for both the classical and quantum reservoir networks. For the training, both networks are executed in an open loop to evolve the reservoir states and calculate the $\pmb{W}_{out}$ matrix. For predictions, both classical and quantum networks evolve dynamically in the closed-loop configuration from ensembles of points in phase space, sampled on the attractor randomly, to quantify the time accuracy. These points correspond to the points in the unseen test data set.
In Tab. 2, we compare the hyperparameters of both quantum and classical reservoirs. Values inside square brackets indicate that the parameters are optimized within this range, multiple values indicate that we repeated the hyperparameter search for each of these values and then selected the best VPT out of them. The random rotation angles ($\alpha$) in the third unitary block can also be treated as a hyperparameter. Alternatively, one could derive them from a uniform random distribution, which we shall do in this paper. For each ensemble, we sample $V(\alpha)\,\,\epsilon\,\,\mathbb{R}^{n}$ from a uniform distribution interval $[0,4\pi]$ with a predefined seed, which we keep it fixed throughout the training and prediction of a particular realization. This is similar to the classical input $\pmb{W}_{in}$ and reservoir weight matrices $\pmb{W}$, which are also pseudo-randomly generated and fixed for any specific realization, as explained in Sec. II.1.
@@ IV.2.1 Time-accurate predictions
In Fig. 8, we present the VPT values derived with classical reservoir computing and compare them with different configurations of quantum reservoir computers from Tab. 1. The error bars at each value indicate the variation arising from different random seeds. We have tested five different seeds for each value. These seeds correspond to randomly generated weight matrices $\pmb{W}$ and $\pmb{W}_{in}$ for the classical reservoir and variation of ${V(\alpha)}$ values for quantum reservoir architectures. For each of these choices, after training, we initialize the closed-loop predictions from 20 different initial points for quantifying VPT and its mean value.
Figure 8: Lorenz-63 system. VPT vs reservoir sizes for both classical and quantum reservoir configurations.
Figure 9: Lorenz-63 system. VPT/circuit depth vs. reservoir sizes for different quantum reservoir configurations.
@@
The prediction performance of the two RF-QRC configurations, QRC-C3 and QRC-C5, do not improve with the reservoir sizes. The QRC-C4 architecture, however, shows comparable performance to the reservoirs with feedback loops (QRC-C1 and QRC-C2). These results highlight the importance of carefully chosen feature maps for QRC.
Figure 10: Lorenz-63 system. Closed-loop time-series predictions. Comparison of true predictions with classical reservoir computing (CRC) and emulated quantum reservoir computing (QRC).
In Fig. 9, the VPT values normalized by the circuit depth are compared between different QRC architectures. We find that the proposed QRC-C4 architecture outperforms all other architectures by demonstrating the best prediction capabilities while requiring a smaller circuit depth.